WENDY'S AND DANIEL'S AGES COMBINED ARE TWICE JAMIE'S AGE. WENDY IS 8 YEARS OLDER THAN DANIEL. JAMIE'S AGE PLUS DANIEL'S AGE IS 20 YEARS. HOW OLD ARE EACH OF THEM?: WENDY'S AND DANIEL'S AGES COMBINED ARE TWICE JAMIE'S AGE. WENDY IS 8 YEARS OLDER THAN DANIEL. JAMIE'S AGE PLUS DANIEL'S AGE IS 20 YEARS. HOW OLD ARE EACH OF THEM?
Answer by stanbon(18792)

WENDY'S AND DANIEL'S AGES COMBINED ARE TWICE JAMIE'S AGE. WENDY IS 8 YEARS OLDER THAN DANIEL. JAMIE'S AGE PLUS DANIEL'S AGE IS 20 YEARS. HOW OLD ARE EACH OF THEM?
-------------------------------------
Let Daniel's age be "x".
Wendy's age is "x+8".
Jamie's age is "x+8 + x" = "2x+8"
----------------------------------
EQUATION:
Jamie's age Plus Daniel's age is 20 years.
2x+8 + x = 20
3x = 12
-----------------
x = 4 (Daniel's age Now)
x+8 = 12 (Wendy's age NOW)
2x+8 = 2*4+8 = 16 (Jamie's age NOW)
========================================
Cheers,
Stan H.